# Mining-Gym: A Configurable RL Benchmarking Environment for Truck Dispatch Scheduling

### v1

@@ IV-B Other Visualization Features
    Cost per Ton ($CPT$): Refers to the financial expenditure incurred for every unit weight (typically one ton) of material produced, processed, or transported.
    $CPT$ $=\frac{C_{KW}+C_{EST}}{P_{VOL}}$
    where $C_{KW}$ = Total known cost and $C_{EST}$ = Total estimated cost
@@
    Fuel Consumption ($FC$): measures the amount of fuel consumed by mining equipment or vehicles relative to the total material hauled during operations:
    $FC=\frac{\text{Total fuel consumed}}{P_{\text{Vol}}}$
    where the Total fuel $\text{Total fuel consumed}=N\times F_{\text{Unit}}$, and the $F_{\text{Unit}}=F_{\text{Unit}}=F_{W}+F_{O}+F_{E}.$
    Here, $N$ is the number of trips completed by all trucks, $F_{W}$, $F_{O}$, and $F_{E}$ are the fuel consumptions for waste, ore, and empty trips, respectively, and $P_{\text{Vol}}$ is the target production volume.
We have also provided an interactive dashboard as part of this benchmark to analyze change in these KPIs in real time.
V Experimental Setup
V-1 Environment Design
We consider a medium sized Bauxite ($5-10$ Million Tons output) mine with 3 Shovels, 20 Trucks, 4 Crushers and 2 Dumping sites, detailed specifications are presented in Table: IV and V. A conceptual diagram of the minesite is provide in Fig. 4.
@@ V-1 Environment Design
TABLE IV: Equipment List with Specifications
V-2 Training Setup
To ensure that the trained agent generalizes well across the testing scenarios, we not only incorporate stochastic parameters to randomize breakdown and repair times for mining equipment and trucks but also enhance the reward function by introducing queue buildup metrics (queue length) and diversity score to prevent repeatedly selecting same shovels.
The parameter values used during training is provided in Table V. For training an RL policy we have used the PPO [29]algorithm from StableBaselines3 [25]). We have used all the default parameter and training settings as present in the repository’s PPO implementation. Note the scenario parameters which help specify the failure scenarios to be simulated during training or testing. For this proof-of-concept implementation and demonstration the results of only $400$ episodes or shifts of training.
Figure 5: Comparison of “Trips per hour” over the shift under different failure scenarios: (a) No failure, (b–d) $(SH_{0},TR_{0})$ values where $SH_{0}$ (Num. Shovels failed) and $TR_{0}$ (Num. Trucks failed) are set as follows—(b) $(0,6)$, (c) $(1,0)$, (d) $(1,6)$. Solid lines represent mean values, and shaded regions indicate variance from repeated runs. For shovels, failures initiate between 100–300 minutes ( 2–5 hours), and for trucks, between 100–500 minutes ( 2–8 hours) within a 12-hour shift.
V-3 Testing Setup
For a random scheduler, the dispatcher’s action $A_{d}$ is selected randomly from the available set of shovel IDs $\{1,2,\dots,SH\}$. The selection follows a uniform distribution,
$P(A_{d}=a|S_{d}=s)=\frac{1}{SH}\quad\text{for all }a\in\{1,2,\dots,SH\}$
i.e. any state $S_{d}=s$, each action $A_{d}=a$ (where $a$ is a shovel ID) is equally likely to be chosen, with a probability of $\frac{1}{SH}$. In contrast, the RL policy $\pi_{\theta}(a|s)$ is a learned policy, where the action $A_{d}=a$ is selected based on the current state $S_{d}=s$ and the parameters $\theta$. The policy is defined as:
$\pi_{\theta}(a|s)=P(A_{d}=a|S_{d}=s)$
This probability is learned through RL algorithm where the agent adjusts $\theta$ based on rewards from its actions, optimizing its decision-making process over time. We tested the schedulers under different challenging scenarios:
  1. 1.
    All equipment available, represented as $(SH_{0},TR_{0})=(0,0)$, i.e., no Shovels and Trucks offline.
  2. 2.
    Six Trucks under maintenance, i.e., $(SH_{0},TR_{0})=(0,6)$.
  3. 3.
    One Shovel offline, i.e., $(SH_{0},TR_{0})=(1,0)$.
  4. 4.
    One Shovel and Six Trucks unavailable, i.e., $(SH_{0},TR_{0})=(1,6)$.
All failure scenarios are repeated according to a consistent temporal schedule over five repeat trials, with the inherent stochasticity in event timings due to probabilistic distributions. These repeats are then considered in the results section for analysis and discussion.
Stochastic Parameters Probability Distributions
Fuel consumption Liters
Loaded Truck (FL) in lt/ton-km $\text{Normal}(0.25,0.001)$
Empty Truck (FE) $\text{Normal}(0.35,0.001)$
Shovel (Lt/hr) $\text{Uniform}(250,300)$
Equipment specifications Value
Truck payload (TP) $\text{Normal}(98,0.5)$ t
Shovel bucket capacity (SCP) $\text{Normal}(43,0.5)$ t
Empty Truck Speed (TS) $\text{Normal}(40,60)$ km/hr
Loaded Truck Speed (TL) $\text{Normal}(25,40)$ km/hr
Equipment Maintenance** Time (in mins )
MTBF Shovel (FSH) $\text{Poisson}(2520)$
MTBF Truck (FTR) $\text{Poisson}(2160)$
MTBF Crusher (FCR) $\text{Poisson}(30000)$
MTBF Dumping Site (FDS) $\text{Poisson}(60000)$
MTTR Shovel (RSH) $\text{Poisson}(240)$
MTTR Truck (RTR) $\text{Poisson}(300)$
MTTR Crusher (RCR) $\text{Poisson}(240)$
MTTR Dumping Site (RDS) $\text{Poisson}(120)$
Loading and Maneuver Time Time (in mins.)
Time Truck Loading (TRL) $\text{Normal}(8,2)$
Travel Time Shovel to Dump (STD) $\text{Normal}(15,5)$
Time Truck Dumping (TRDM) $\text{Normal}(5,1)$
Travel Time Dump to Shovel (DTS) $\text{Normal}(15,5)$
Travel Time Shovel to Crusher (STC) $\text{Normal}(15,2)$
Time Truck Unloading at Crusher (TRCR) $\text{Normal}(5,1)$
Travel Time Crusher to Shovel (CTS) $\text{Normal}(15,2)$
Scenario Parameters -
Shovel To Fail (STF) Binomial (2, 0.5)
Trucks To Fail (TTF) Binomial (6, 0.5)
Shovel Initial Breakdown (SIB) Uniform (100,300)
Truck Initial Breakdown (TIB) Uniform (100,500)
@@ V-3 Testing Setup
TABLE V: Distributions for stochastic parameters used in the example. MTBF: Mean Time Between Failures, MTTR: Mean Time To Repair.
Figure 6: Comparison of production volume $P_{vol}$ in metric tonnes per shift, averaged over 5 repeats trials.
VI Results and Discussion
The RL scheduler demonstrated significant performance improvements over the random scheduler across all scenarios, see Fig.5 and Fig. 6 :
  • Trips Per Hour: RL increased productivity by 18.2% on average, with the greatest improvement (27.2%) observed in the most challenging scenario (1 shovel and 6 trucks offline).
  • Production Volume: RL delivered 19.7% higher production volumes on average, with improvements ranging from 14.3% to 23.3% across all scenarios.
The performance advantage of RL was most pronounced in scenarios with multiple resource constraints, suggesting that RL’s optimization capabilities become more valuable as operational complexity increases. As is showing through the gain in production volume (see Fig.6) RL improved production volume across different scenarios. In the (0,0) case, where all equipment was available, the increase was 18.4%. With six trucks offline (0,6), the improvement rose to 22.7%. When one shovel was offline (1,0), production volume increased by 14.3%. Under both constraints (1,6), RL achieved the highest improvement at 23.3%.
The RL scheduling (RL-PPO) consistently outperforms random scheduling baseline method across multiple metrics and scenarios, maintaining higher productivity, especially mid-shift. Its intelligent decisions optimize resource use, making it highly effective in real-world mining operations with equipment constraints, leading to increased production.
VII Conclusion and Future Work
The introduction of Mining-Gym addresses a critical gap in mining process optimization by providing a configurable, open-source benchmarking environment for RL algorithm evaluation in truck dispatch scheduling. Our framework offers high-fidelity DES modeling of real-world complexities, seamless integration with OpenAI Gym, and comprehensive visualization tools. Preliminary experiments show that RL-based scheduling achieves up to 27.2% higher productivity and 23.3% greater production volume, with the most significant gains in resource-constrained scenarios. Mining-Gym provides a standardized evaluation framework, ensuring reproducible research and supporting the adoption of RL solutions in industrial mining operations.
The development of Mining-Gym as a benchmarking environment will continue in the following directions:
  • Enhanced Simulation Fidelity: Future iterations will improve traffic modeling, incorporate geographical factors, model heterogeneous equipment fleets, and include additional operational constraints to enhance simulation realism.
  • Expanded Scenario Library: We will develop a broader set of standardized test scenarios to cover diverse challenges such as weather disruptions, varying ore characteristics, and dynamic production targets.
  • Multi-objective Optimization Benchmarks: The framework will be extended to support multi-objective optimization, balancing production, cost, equipment lifespan, and environmental impacts—reflecting real-world mining decision-making.
  • Integration with Digital Twin Capabilities: Enhancing Mining-Gym with digital twin functionalities will connect simulations with real-world data, facilitating data-driven optimizations and bridging the gap between simulation and operational implementation.
Acknowledgments

### v2

@@ Appendix A Modeling Assumptions and Abstractions
    Trailing transport: Truck cycles transporting tailings to the dumping site are not considered in the current version of the simulator, as they have lower priority and often use separate equipment. Including tailings would add unnecessary complexity without significantly improving the efficiency of primary ore transport operations.
@@
    Environment Reset: During RL-based training, the environment operates in shifts (episodes). Each shift begins with trucks scheduled to shovels using the Equal-Queue scheduler by default. The shift ends when the time limit is reached, ignoring trucks still in transit. Only completed trips and transported material within the shift’s time frame are counted.
Appendix B Experimental Scenario Parameters
Table 7: Equipment List with Specifications
@@ Appendix B Experimental Scenario Parameters
Table 8: Distributions for stochastic parameters. Parameters are grouped as Sample Once (fixed per environment) and Sample per Episode/Shift (varies daily). MTBF: Mean Time Between failures, MTTR: Mean Time To Repair.
Parameter Group Parameter Distribution / Value
A. Sample Once (Environment-fixed)
Equipment counts Trucks (TR) 60
Shovels (SH) 8
Crushers (CR) 5
Dumps (DS) 3
Equipment characteristics Truck payload (LO) 100 t
Empty Truck Speed (TE) Normal(55, 12) km/hr
Loaded Truck Speed (TL) Normal(30, 8) km/hr
Shovel bucket capacity (SCP) Not specified
Equipment Maintenance MTBF Shovel (FSH) Poisson(75)
MTBF Truck (FTR) Poisson(90)
MTBF Crusher (FCR) Poisson(120)
MTBF Dumping Site (FDS) Poisson(180)
MTTR Shovel (RSH) Poisson(60)
MTTR Truck (RTR) Poisson(45)
MTTR Crusher (RCR) Poisson(90)
MTTR Dumping Site (RDS) Poisson(30)
Infrastructure travel times Shovel → Dump (SZxD) Zone 1: Normal(18, 3)
Zone 2: Normal(12, 2), Zone 3: Normal(8, 1)
Shovel → Crusher (SZxC) Zone 1: Normal(8, 2)
Zone 2: Normal(15, 3), Zone 3: Normal(22, 4)
B. Sample per Episode / Shift (Varies Daily)
Fuel consumption Truck with ore (FO) Normal(0.28, 0.05) lt/ton-km
Truck empty (FE) Normal(0.20, 0.06) lt/ton-km
Shovel (epsilon) Normal(0.35, 0.07) lt/hr
Truck Loading Times Zone 1 (TRL_C1) Normal(6, 1) min
Zone 2 (TRL_C2) Normal(9, 2) min
Zone 3 (TRL_C3) Normal(12, 2) min
Truck Dumping / Unloading At Dump (TRDM) Normal(5, 2) min
At Crusher (TRCR) Normal(15, 1) min
Experimental Variations Shovels to fail (STF) Uniform(1, 5)
Trucks to fail (TTF) Uniform(10, 30)
Shovel initial breakdown (SIB) Uniform(10, 45)
Truck initial breakdown (TIB) Uniform(10, 90)
Table 9: Reward function hyperparameters used in demonstration experiments, balancing production efficiency ($\omega_{1}$, $\omega_{2}$), performance bonuses and thresholds ($b_{1}$, $b_{2}$, $\theta_{1}$, $\theta_{2}$, $\phi_{1}$, $\phi_{2}$), and immediate reward parameters ($\alpha$, $\beta$, $\gamma$, $\delta$, $k$, $\lambda$, $|W_{\text{div}}|$, $|W_{\text{streak}}|$). All parameters are configurable in Mining-Gym for alternative objectives and ablation studies.
Function Component Parameter Description Value
Episodic Reward (Eq. 5) $\omega_{1}$: Production efficiency weight 0.4
$\omega_{2}$: Diversity penalty weight 0.6
Immediate Reward Weights (Eq. 3) $\alpha$: Trip time penalty weight 0.25
$\beta$: Shovel queue time penalty weight 0.15
$\gamma$: Diversity score penalty weight 0.30
$\delta$: Streak penalty weight 0.30
Performance Bonus (Eq. 6) $b_{1}$: High-performance bonus value 0.5
$b_{2}$: Moderate-performance bonus value 0.2
Production Thresholds (Eq. 6) $\theta_{1}$: High production threshold 0.95
$\theta_{2}$: Moderate production threshold 0.90
Diversity Thresholds (Eq. 6) $\phi_{1}$: High diversity threshold 0.65
$\phi_{2}$: Moderate diversity threshold 0.50
Exponentially Weighted Sliding Window (Eq. 4) $k$: Window size (decision points) 5
$\mu$: Exponential decay rate 0.5
$n$: Effective window size $\min(k,|D|)$
Diversity Score Calculation $|W_{\text{div}}|$: Recent decisions window 30
$\rho$: Numerical stability constant $10^{-10}$
Streak Penalty Calculation $|W_{\text{streak}}|$: Consecutive decision window 15
Table 10: Truck Specifications
Maker Model Payload (MT) Speed (loaded km/hr) Fuel Cons. ($t\textperiodcentered km^{-1}$)
Komatsu 830E-2 231 59-64 0.3 - 0.4 liters
CAT 793F 231 55-60 0.35 - 0.45 liters
Terex MT4400 AC 218 60-64 0.35 - 0.45 liters
CAT 777(05) 98.2 62 - 67.1 0.25 – 0.35 liters
Table 11: Shovel and Crusher Specifications
Model Fuel per Hour Running Cost (USD)
Shovel Specifications
CAT 6040/6050 Hydraulic Shovels 250-300 liters per hour \$3.50 to \$4.50 /t
Komatsu PC8000 946 liters per hour \$4.00 - \$5.00 / t
Hitachi EX8000-6 1100 liters per hour \$4.50 - \$5.50 / t
Crusher Specifications
Metso Lokotrack LT106 1.5 to 2 gallons of diesel per hour \$4.50 to \$6 per hour
Appendix C Event-driven or decision-point-based RL setting
The DES mining environment requests decisions from the RL policy only at specific decision points, such as when a truck requires a shovel assignment. This event-driven approach, illustrated in Fig.6, ensures efficient agent-environment interaction aligned with real mining operations, where continuous interaction is impractical. The agent observes the current state (including shovel and truck attributes) following the Load-Haul-Dump-Return-Query (LHDRQ) cycle, and makes resource allocation decisions using learned dispatching policies. As shown in the figure, the RL policy processes environment states to optimize shovel assignments based on multiple factors, including queue length and equipment status. The simulation handles stochastic elements like equipment failures through the Preemption Handler, while the Resource Handler manages resource availability and assignments.
Figure 6: The DES mining environment requests decisions from the RL policy only at specific decision points, such as when a truck requires a shovel assignment.